Weather: Progress report for the month of February

Perhaps, this has been the month with the most progress. In this month I started a project to make a Java applet which will make isosurfaces and display this into VRML. I also met with the CAPS personnel. They will originally supply me with an initial data set of the Syracuse region, during the May 1 time period. I have also set up another meeting with Bob Farallel for latter this month, although no exact date has been set.

At the beginning of this month, Geoffrey and I discussed the plans to develop a set of Java applets which will read in a three dimensional scientific data set, probably in HDF format, apply a marching cubes algorithm at a user chosen isosurface value, and then output this into a VRML rendering window. The marching cubes algorithm, is a one taken from the Lorensen and Cline algorithm. I have an old c program and will convert this into Java. Since I am new to Java, my first goal was to learn the basics of this language. After that, I developed a stand alone Java program do solve Poission’s equation in two-dimensions implementing a Gauss Seidel iterative solver. Currently, my plans are to convert this program into a Java applet, where a widget will be introduced, which will define an overelaxation parameter. The applet would then include the capabilities to plot the solution. The next task would then be to turn the old marching cubes program into Java.

I met with several key personnel at the Center for Analysis and Prediction of Storms, (CAPS), on April 30, 1996. The first person I met with was David Jahn who is the assistant director of the center. He proceeded to give me a standard overview of the center, and became my tour guide during the day. He explained to that the no one has ever tried to use the code to model winter time scenarioΦ. He also explained to me that the center has a very limited staff, and it would be doubtful if they could help me much. Another point that David mentioned was that state of Oklahoma has a very advanced weather prediction system because they have a advanced Mesonet system. Their Mesonet system is where at each country, they have radar systems set up to read in the weather conditions. The data is then transferred to their place, and analyzed. Their data can be obtained from http://wwwcaps.gcn.uoknor.edu/Weather.html. This initial meeting went on for two hours.

After the meeting with David, I was taken for a meeting with Howard Johnson, who is head of the Oklahoma Mesonet center. Again, he quickly pointed out who sophisticated the Oklahoma setup was in comparison to other states. To see the difference between Oklahoma and it’s surrounding states, look at the web page from the proceeding paragraph, and look under the Oklahoma Mesonet/NWS station plots. Here we can clearly see the concentration of data around Oklahoma in comparison to Texas. In other words, NY is much like TX such that the gathering of initial data will be much less accurate than that of Oklahoma. This will then greatly effect the accuracy of the weather prediction from their code.

Next I had a working lunch with Kelvin Droegemeier, the director of the center. The lunch went extremely well, and Kelvin and I discussed several issues. The most important issue is that Kelvin told me that he wants this project to happen for several reason, but the only problem is their lack of resources. He told me that he would talk to a friend of his at SUNY, to see if they could also help us. He also became extremely interested in our parallel implementation of Adaptive Mesh Refinement, (AMR). He quickly pointed out to me that their parallel code does not work with AMR and this is a severe limitation of their code. Hopefully interest in this might help him put some of their effort into winter modeling around the Syracuse region.

After this meeting we meet with Paul Janish who is in the NSSL (National Severe weather group). Paul discussed with me the efforts of the people at Buffalo, and their efforts to predict lake effect snow. He mentioned that we should try to get in touch with their (MM5) group. He also stated that if he does get in touch with them for us, he will email me to give me some contact names.

The next meeting was with Fred Carr, who is head of the winter time modeling group. Fred told me about their current efforts to model snow in the Oklahoma region. He has a graduate student, John Case (who is from the Rochester region), trying to do this. I asked him about the possibility of this student trying to model lake effect snow in the Syracuse region, and I was told a few interesting facts. The first fact is that this student actually wants to do this. The next fact is that this student is trying to work on a masters thesis, and it would be doubtful if this student would change his plans since he has already started. The fact is once again the difficulty to predict the weather accurately here, since the Mesonet is not advanced compared to the Oklahoma region. They are very worried about their code being used to predict the weather inaccurately since they feel that this could politically harm their efforts. Fred told me that he would have a meeting with Kelvin to possibly talk about letting this student, or another student, model the Syracuse region. I was also informed that if we supported a new graduate student of theirs, then after about 1 year of work, they would be able to get a decent prediction of lake effect snow. In a few months the code should be able to predict the snow, although it would be inaccurate.

The last meeting was with several researchers including Ming Xue (head of the code development), Gene Bassett (head of the visualization support), Steve Weygandt (a graduate student), Keith Brewster (a research scientist), and once again David Jahn. Although the people at the meeting were quickly saying that this project would be interesting and they would want to help us, David Jahn quickly pointed out that they were not allowed. The outcome of this meeting was good though. Keith Brewster will supply me with an initial data set of the Syracuse region. Gene Bassett will work on this data such that we will be able to see some cloud formation (typically clouds only form in the code after running the code for several hours). Gene will also help me with other visualization problems. Ming will try to evolve this initial data, some time in June, but the code will be running inaccurately. Possibly by the end of the summer the code will be predicting lake effect snow, although the prediction will be very inaccurate in space. The researchers also informed me that it would be impossible for me to run their code, unless they have set up the parameters and run it.